You are an ice cream maker and decided ice cream w/tih 9% butterfat is the best. How many gallons of ice cream with 5% butterfat should be mixed with ice cream that is 15% butterfat to make 200 gallons of desired blend?
Answer by macston(5194)

X=gallons of 5%; Y=gallons of 15%
X+Y=200gal
X=200gal-Y
0.05X+0.15Y=0.09(200) Substitute for X
0.05(200gal-Y)+0.15Y=18gal
10gal-0.05Y+0.15Y=18gal Subtract 10gal from each side.
0.10Y=8gal
Y=80 gallons ANSWER 1: There should be 80 gallons of 15%.
X=200gal-Y=200gal-80gal=120gal ANSWER 2: There should be 120 gallons of 5%.
CHECK:
0.05X+0.15Y=0.09(200)
0.05(120gal)+0.15(80gal)=18gal
6gal+12gal=18gal
18gal=18gal